OpenStudy (anonymous):

[okay the full prob is 2x-3/5x+1 Multiplied by 15x^2-7x-2/6x^2-13x+6

OpenStudy (anonymous):

(2x-3/5x+1)(15x^2-7x-2/6x^2-13x+6) nominator = (2x - 3)(15x^2 - 7x - 2) = 30x^3 - 14x^2 - 4x - 45x^2 + 21x + 6 denominator = (5x + 1) (6x^2 - 13x + 6) = 30x^3 - 65x^2 + 30x + 6x^2 - 13x + 6 i hope u can simplify this

OpenStudy (anonymous):

uhm that is just too confusing and i dont think its right. i really just need to know if 6x^2-13x+6 can be factored

OpenStudy (anonymous):

(2 x-3) (3 x-2)

OpenStudy (shadowfiend):

Suzi's right!